Consider an economy having following values of Consumption, Investment, Government

 

Spending, and Taxes, 

 

C = 300 + 0.7Yd         I = 200 – 200i G = 600          T = 100   

 

Derive the aggregate demand equation. Also, calculate the output in an economy.  


1
Expert's answer
2021-06-02T12:26:31-0400

Aggregate demand shows the total demand for domestically produced goods and services by domestic residents and foreigners. It is the sum of consumption expenditure, investment expenditure and government expenditure.

"AD = C+ I + G"

"= 300 + 0.7Yd + 200 - 200i + 600 \\

 = 1100 + 0.7(Y -T) - 200i \\

 = 1100 + 0.7Y - 0.7(100) -200i \\

 = 1100 + 0.7Y -70 -200i \\

 = 1030 + 0.7Y - 200i \\"

 

The market is in equilibrium when real output is equal to the aggregate demand. Real output is the market value of final goods and services produced in the economy. 

 "Y = AD \\

 Y = 1030 + 0.7Y - 200i \\

Y - 0.7Y = 1030 -200i \\

 0.3Y = 1030 -200i \\

 Y = 3433.3 -666.7i \\"